OpenStudy (anonymous):

what is the simplified form of x2-16/x+4

OpenStudy (anonymous):

Is it \[x^2~~~?\]

OpenStudy (anonymous):

\[x-4\]

OpenStudy (anonymous):

because if it is, the answer is x-4 because x^2-16=(x+4)(x-4) The (x+4)s cancel out, and what's left is x-4

OpenStudy (anonymous):

\[\frac{x^2-16}{x+4}=\frac{(x+4)(x-4)}{x+4}=x-4\]

OpenStudy (anonymous):

yes it is x^2

Directrix (directrix):

{(x + 4)(x - 4)] / (x+4) = x - 4 where x cannot be 4